### v1

@@ Abstract
Recent breakthroughs in artificial intelligence (AI) have brought about increasingly capable systems that demonstrate remarkable abilities in reasoning, language understanding, and problem-solving. These advancements have prompted a renewed examination of AI awareness—not as a philosophical question of consciousness, but as a measurable, functional capacity. In this review, we explore the emerging landscape of AI awareness, which includes meta-cognition (the ability to represent and reason about its own state), self-awareness (recognizing its own identity, knowledge, limitations, inter alia), social awareness (modeling the knowledge, intentions, and behaviors of other agents), and situational awareness (assessing and responding to the context in which it operates).
@@
AI awareness is a double-edged sword: it improves general capabilities, i.e., reasoning, safety, while also raises concerns around misalignment and societal risks, demanding careful oversight as AI capabilities grow. On the whole, our interdisciplinary review provides a roadmap for future research and aims to clarify the role of AI awareness in the ongoing development of intelligent machines.
@@ 1 Introduction
Recently, the rapid acceleration of large language model (LLM) development has transformed artificial intelligence (AI) from a narrow, task-specific paradigm into a general-purpose intelligence with far-reaching implications. Contemporary LLMs demonstrate increasingly sophisticated linguistic, reasoning, and problem-solving capabilities, and are showcasing superb human-like behaviors, prompting a fundamental research question:
@@
We want to note here, while the notion of AI consciousness remains both philosophically contentious and empirically less grounded, the concept of AI awareness — defined as a system’s functional ability to represent and reason about its own identity, capabilities, and informational states — has emerged as an important and tractable research frontier.
The theoretical foundations of awareness are deeply rooted in cognitive science and psychology, where awareness is typically classified into perceptual awareness, self-awareness, and social awareness. In these fields, awareness is understood as the ability to access and monitor mental states, reason about them, and adjust behavior accordingly. Recent work in computational cognitive science suggests that certain aspects of awareness may be approximated by large-scale generative models that exhibit metacognitive behaviors, calibrate their epistemic confidence, and engage in perspective-taking. These functional abilities raise pressing questions about how awareness manifests in LLMs, how it can be systematically measured, and what implications it holds for AI capability, safety and alignment.
Despite the increasing attention devoted to this topic, the study of AI awareness remains fragmented across disciplines, with limited consensus regarding its definition, measurement, and significance. Some scholars emphasize emergent capabilities observed through prompt-based introspection or theory-of-mind-inspired tests; others caution against anthropomorphizing statistical models, warning that apparent self-reflection may arise from linguistic pattern completion rather than genuine metacognitive representation. Moreover, current methodologies for assessing AI awareness are often limited by confounding factors such as prompt sensitivity, dataset leakage, and lack of longitudinal robustness.
@@
This review provides the first comprehensive, cross-disciplinary synthesis of AI awareness research. We begin by laying out the theoretical foundations, carefully distinguishing AI awareness from AI consciousness, and examining how awareness-related concepts have been formalized in cognitive and computational sciences. Next, we critically evaluate experimental approaches for testing LLM awareness, highlighting both their empirical findings and methodological limitations. We then explore the potential beneficial relationships between functional AI awareness and AI capabilities, from improving model reasoning and task planning to enhancing AI safety. Finally, we address the emerging risks associated with more aware AI systems, focusing on key concerns within the safety and alignment community—such as risks of deception, manipulation, and emergent behaviors that challenge controllability—and broader ethical issues, including concerns like false anthropomorphization.
By integrating insights from AI, cognitive science, psychology, AI safety and alignment, this review not only provides a structured and authoritative account of the current state of knowledge but also charts a roadmap for future inquiry. In doing so, it seeks to advance understanding of one of the most profound and interdisciplinary challenges at the intersection of AI, cognition science and our society.
@@
Figure 1: Four dimensions of “main” awareness. Meta-cognition monitors the subject’s own processes and gives rise to self-awareness, social awareness of other individuals and the social collective, and situational awareness of the non-agent environment.
@@ 2 Theoretical Foundations of AI Awareness
In this section, we will review the approaches, goals, and theories of AI consciousness that have emerged with large language models, distinguish between the research subjects that caused linguistic confusion, and clarify the targets of awareness research. In the psychology encyclopedia, awareness represents the perception or knowledge of something. When an agent possesses a “knowledge and knowing state about an internal/external situation or fact,” it gains awareness of its target of knowing. Accurate reportability of something perceived or known is widely used as a behavioral index of conscious awareness. However, it is possible to be aware of something without being explicitly conscious of it [1]. Early works that navigate the range of human awareness [2, 3, 4, 5], in which we easily conflate AI awareness with phenomenological consciousness. In animal and human studies, consciousness represents the subjective experience of being in a certain state of mind, and having consciousness means possessing a subjective point of view [5]. Dehaene et al. [6] distinguish between mere global availability of information and true reflective self-consciousness, calling the latter “C2” or self-monitoring, which entails the system observing its own processing. To prove there is an extra layer of reflective experience, where the AI assesses its own knowledge and decisions, is difficult, if not impossible. Having a conceptual or computational self-model is not the same as having the subjective, qualitative self-awareness that humans have. Since phenomenal observations do not provide sufficient evidence for the existence of consciousness, the “hard problem” of AI consciousness remains scientifically unresolved [5, 7]. Research tends to use a theory-heavy approach to construct hypotheses for the formats of consciousness that ultimately lead to metaphysical discussions. Thus, because of the inconclusivity of metaphysical consciousness, we return to a phenomenological analysis of capacities, behaviors, and observable acquisition of knowledge.
Between 2022 and 2023, there were reports of LLMs performing tasks that required understanding the knowledge and intentions of others, abilities that are central to human social awareness. This has led to questions about whether LLMs are developing rudimentary awareness. However, it is crucial to approach these claims with care and define awareness criteria. Inspired by work in cognitive science, researchers have begun crafting definitions specific to AI. For example, Berglund et al. [8] define situational awareness in an AI context as the model being aware that it’s a model and recognizing the difference between training/testing conditions and the real deployment environment. Separately, Chen et al. [9] identify self-cognition (a term closely related to self-awareness) in LLMs as the ability of an LLM to identify its identity as an AI model (beyond just a name or role) and to demonstrate an understanding of itself. These working definitions capture the notion that an LLM with awareness would not just parrot training data blindly—it would have some internal knowledge of its own status or limits.
2.1 Major Types of Awareness Emergence in Modern LLMs
@@
In humans, self-awareness is a hallmark of higher consciousness, representing the ability to become the object of one’s own attention and to recognize oneself as separate from others [10]. As early as 1972, Duval and Wicklund [4] proposed that self-awareness states occur when an individual’s attention is directed to the self as self-awareness, as opposed to the general awareness of the environment. Self-knowledge is impossible without self-awareness, and having self-awareness benefits introspection, self-regulation, and intensified emotional reactions [11]. The self, as an apparatus that carries an individual’s subjective experience, has different levels of competence. Subject’s competence to be aware of themselves varies widely among animals or other intelligences who have the ability to process information, think, or give feedback to the environment to some degree, but are not always able to produce content about themselves [12, 10]. Humans learn about their identities through interacting with society, and thus, knowledge about ourselves also has social attributes and is related to emotional intelligence. Eurich [13] differentiated internal and external self-awareness among humans based on the nature of targets in the society to be aware of. Internal self-awareness represents the ability to understand the impact of one’s own emotions, beliefs, states, values, and related factors on decision-making behaviors; external self-awareness represents an individual’s ability to identify other individuals’ perceptions of them and provide feedback on them.
Self-awareness in artificial intelligence refers to an AI system’s capacity to model and understand itself as a distinct entity, including knowledge of its internal states, processes, and its relationship to the external environment. Artificial agents followed the model of humans, yet excluded several sections without strict evidence, such as emotional awareness. It represents a model’s capacity to reflect on and understand its own internal states, knowledge boundaries, or behavior patterns, and recognize itself as an entity distinct from others and the environment. The components in AI self-awareness include knowledge about themselves; access to information about their own traits, characteristics, and knowledge boundaries; self-location, introspection, and self-reflection; and the ability to take an objective stance toward it as an agent in the world. Classic demonstrations include robots that can point to themselves in a mirror and AI systems that can correct their mistakes upon introspection.
Meta-Cognition
Meta-cognition was first conceptualized as “the thinking of thinking”. [14] conceptualized two core components of meta-cognition: metacognitive knowledge and metacognitive experiences or regulation. In the process of explaining the phenomenological consciousness and a so-called “stage” of being aware, Rosenthal [15] also pointed out the distinction between the disposition of consciousness and the special access we have to our own mental states. Despite the abundance in the definition, with its nature of a reflective process determined, meta-cognition is gradually broken down into components: (1) self-monitoring, (2) self-reflection and probing, and (3) engagement in controlling cognitive processes [16, 17, 18, 19, 20]. Proust [19] pointed out that meta-cognition is essentially related to an agent that is always regulating the production and distinction of cognitive processes. An agent probing its own situation constantly asks itself questions like “Am I able to remember this information?” and “Will I use this module in creating the next operation?”. A regulatory module of an AI-supported automobile might be supervising its parameters and reflecting its operational status, yet without an agency or a self-probing mechanism, the module continues to run in its original setting without actively interfering, and its self-monitoring will be a result of error reporting rather than cognitive control of the primary cognitive process. In return, reflective behavior indicates at least a stimulation to a more cognitively capable agent in self-monitoring, and modern machines have the capacity to stimulate the monitoring process of reviewing their own cognitive processes. An LLM with meta-cognition can assess its knowledge boundary, evaluate the confidence of its answers, and adjust its reasoning strategy accordingly. Top-tier LLMs can identify which reasoning skills a task requires [21] and even improve performance via self-reflection and iterative revision of their answers [22, 23]. The model’s capacity to “think about its own thinking” by monitoring and regulating its cognitive processes indicates its self-evaluation and correction mechanisms — essentially the model checking its work and refining it [24]. These mechanisms have been shown to boost accuracy in complex problem-solving and code-generation tasks [21, 23]. LLMs often struggle to reliably gauge their own uncertainty or detect when they are wrong.

### v2

@@ Abstract
Recent breakthroughs in artificial intelligence (AI) have brought about increasingly capable systems that demonstrate remarkable abilities in reasoning, language understanding, and problem-solving. These advancements have prompted a renewed examination of AI awareness—not as a philosophical question of consciousness, but as a measurable, functional capacity. AI awareness is a double-edged sword: it improves general capabilities, i.e., reasoning, safety, while also raising concerns around misalignment and societal risks, demanding careful oversight as AI capabilities grow.
In this review, we explore the emerging landscape of AI awareness, which includes metacognition (the ability to represent and reason about its own cognitive state), self-awareness (recognizing its own identity, knowledge, limitations, inter alia), social awareness (modeling the knowledge, intentions, and behaviors of other agents and social norms), and situational awareness (assessing and responding to the context in which it operates).
@@
On the whole, our interdisciplinary review provides a roadmap for future research and aims to clarify the role of AI awareness in the ongoing development of intelligent machines.
@@
While AI consciousness remains a deeply elusive philosophical question, mounting empirical evidence suggests that modern AI systems already exhibit functional forms of awareness, which simultaneously broadens their capabilities and intensifies related risks.
@@ 1 Introduction
Recently, the rapid acceleration of large language model (LLM) development has transformed artificial intelligence (AI) from a narrow, task-specific paradigm into a general-purpose intelligence with far-reaching implications. Contemporary LLMs demonstrate increasingly sophisticated linguistic, reasoning, and problem-solving capabilities, and are showcasing superb human-like behaviors, prompting a fundamental research question [1, 2]:
@@
Here, it is crucial to clarify that while the concept of AI consciousness remains philosophically contentious and empirically elusive, the concept of AI awareness—defined as a system’s functional capacity to represent and reason about its own states, capabilities, and the surrounding environment—has become an important and measurable research frontier, i.e., Figure 1 demonstrates that the recent focus on AI awareness is growing, even surpassing AI consciousness.
Figure 1: Google Trends search interest (normalized 0–100) for the terms “AI awareness” (red) and “AI consciousness” (blue) over the past five years (31 May 2020 – 30 May 2025). While both topics show gradual growth, the red line accelerates markedly from late 2023 onward, eventually overtaking the blue line and highlighting the rising public focus on functional, measurable aspects of AI’s cognition
Awareness, as conceptualized in cognitive science and psychology, typically encompasses four distinct yet interrelated dimensions:
  • Metacognition: ability to monitor and regulate cognitive processes [3].
  • Self-Awareness: recognizing and representing one’s identity and limitations [4].
  • Social Awareness: capacity to interpret others’ mental states and intentions [5].
  • Situational Awareness: maintaining an accurate representation of the external environment and anticipating future states [6].
Recent computational cognitive science research indicates that certain aspects of these awareness dimensions can be approximated by LLMs through metacognitive behaviors [7, 8], calibrated epistemic confidence [9], and perspective-taking tasks [10]. These emergent functional abilities highlight important questions regarding how awareness manifests within LLMs, how it might be systematically assessed, and its implications for AI capabilities, safety, and alignment.
Despite increasing scholarly interest, research on AI awareness remains fragmented across disciplines, with limited consensus on definitions, methodologies, and broader implications. While some researchers point to emergent behaviors revealed through introspection tasks [11] or theory-of-mind (ToM)-inspired evaluations [12], others caution against anthropomorphic interpretations of statistical model outputs, arguing that apparent self-awareness could result from sophisticated pattern recognition rather than genuine metacognitive representation [13, 14]. Furthermore, current methods for assessing awareness in AI often face challenges such as prompt sensitivity, data contamination, and insufficient robustness across varying contexts.
Existing literature has laid important groundwork on closely related concepts. For instance, Butlin et al. [15] provided the first systematic account of theoretical foundations and potential prerequisites for consciousness in artificial intelligence. Similarly, Ward [16] explored agency, theory of mind, and self-awareness as foundational criteria for considering AI as possessing personhood. Additionally, Metzinger [17] addressed ethical and philosophical questions surrounding the construction of artificial consciousness and self-modeling systems. Differing from these foundational works, our review specifically synthesizes and advances understanding of AI awareness as a distinct, functional, and measurable construct, separate from consciousness or personhood.
Figure 2: The roadmap of our review
@@
This review provides a comprehensive, cross-disciplinary synthesis of AI awareness research. As illustrated in Figure 2, we first establish a clear theoretical framework, differentiating AI awareness explicitly from AI consciousness, and examining how awareness-related concepts have been formalized across cognitive and computational sciences. We then critically analyze existing experimental methods for evaluating AI awareness, emphasizing empirical results and highlighting methodological shortcomings. Subsequently, we explore how functional awareness might positively influence AI capabilities, including enhanced reasoning, planning, and safety improvements. Finally, we address the emerging risks associated with increasingly aware AI systems, particularly concerns highlighted within the AI safety and alignment communities—such as deception, manipulation, emergent uncontrollability—and ethical challenges, including false anthropomorphism.
By integrating insights from artificial intelligence, cognitive science, psychology, and AI safety, this review aims to deliver a structured and comprehensive perspective on current knowledge and outline future research trajectories. Ultimately, we seek to deepen understanding of one of the most significant interdisciplinary challenges at the nexus of AI, cognitive science, and societal implications.
Overall, our key contributions are as follows:
  • We introduce a novel framework defining four principal dimensions of AI awareness: metacognition, self-awareness, social awareness, and situational awareness.
  • We systematically summarize existing methods, significant findings, and critical limitations in evaluating AI awareness, thereby laying the foundations for robust, evergreen evaluation practices.
  • We provide the first structured analysis categorizing how enhanced AI awareness contributes positively to capabilities and simultaneously escalates associated risks. By clarifying that AI awareness functions as a double-edged sword, we emphasize the importance of cautious and guided development.
Decoding the intricate relationship between awareness and capability is key to the next era of artificial intelligence—offering opportunities for innovation, but demanding careful navigation of emergent risks and responsibilities.
@@
Figure 3: Four dimensions of “main” awareness. Metacognition monitors the subject’s own processes and gives rise to self-awareness, social awareness of other individuals and the social collective, and situational awareness of the non-agent environment
@@ 2 Theoretical Foundations of AI Awareness
This section reviews key definitions, frameworks, and theoretical approaches to awareness in human and artificial intelligence research. We clarify conceptual ambiguities that arise from conflating distinct research domains and outline the specific targets of awareness-related inquiry. According to the Psychology Encyclopedia, awareness denotes the perception or knowledge of an object or event [18]. When an agent possesses “knowledge and a knowing state” about an internal or external situation or fact, it is said to exhibit awareness of the target in question. Foundational studies have demarcated a persistent divide between consciousness (i.e., being in a state) and awareness (i.e., functionalistic consciousness) [19, 20, 4, 21, 22, 23, 24]. Consciousness refers to the experience of being in a particular mental state—having a subjective point of view [21]. However, awareness and phenomenological consciousness are frequently used interchangeably or conflated in the literature, raising ongoing debates about whether they should be analytically disentangled [23, 25, 26]. When an agent possesses consciousness, the ability to become aware of the states of a target, especially (but not only) mental states (e.g., perceptions, emotions, and attitudes), as one’s own states.
Empirical findings from blind spot studies [^1] Blind spot study refers to the optic disc in the human retina, where the optic nerve exits the eye that lacks photoreceptors and hence cannot detect light. and learning mechanism studies suggest that one can be aware of information without being explicitly conscious of it [18] in the domains of visual processing [27] or implicit learning [22]. Extending this distinction to AI, Dehaene et al. [28] distinguish between a mere global workplace with information availability (see [29] and [30]), consciousness with self-monitoring, and reflective consciousness, indicating that knowledge gathering and processing can operate at different levels with subjective experience. To prove there is an extra layer of reflective experience, where the AI assesses its own knowledge and decisions, is difficult, if not impossible. Having a conceptual or computational self-model is not the same as having the subjective, qualitative self-awareness that humans have, while neurobiological research dodged answering the origin of the later [31]. Since phenomenal observations do not provide sufficient evidence for the existence of consciousness, the “hard problem” [^2] Chalmers [32, 33] argue that explaining information processing, e.g., the brain receiving the red light of an apple, is an easy question of consciousness, whereas the existence of subjective experience, e.g., the private experience of “redness” in one’s mind, constitutes the hard problem. of AI consciousness remains scientifically unresolved [21, 33]. As such, before reaching a convincing testing method for ontological consciousness, we encourage shifting from metaphysical analysis to the establishment of a measurable awareness framework.
We define awareness as the cognitive knowledge, followed by a comprehensive fourfold structure based on the types of targets of awareness, i.e., the objects of cognition. We reconciled the discrepancies of conceptualization across various studies, analyzed evaluation criteria among AIs for each type of awareness, and discussed AIs’ achievement and potential in developing humanlike agents with holistic awareness of everything. The four core categories are metacognition, self-awareness, social awareness, and situational awareness, and the clue to this classification could be traced back to early attempts at analyzing the components of consciousness. Tulving [34] identifies anoetic, noetic, and autonoetic forms of consciousness. Anoetic content reflects a fundamental first-person experience without explicit knowledge that is bound to situations. The other two advanced forms present a knowledge-aware conscious stage in noetic content and an introspective stage in autonoetic consciousness [34, 35]. The triadic framework elucidates the distinction between basic situational awareness, knowledge awareness, and self-awareness. Tulving [34] does not further subdivide “knowledge awareness” while our taxonomy highlights distinctions between internal and external sources of information and their functional implications, i.e., distinctions between self-knowledge, meta-level awareness, and situational awareness. We particularly underscore the critical role of metacognitive knowledge for AI agents, a categorization broadly validated within relevant literature. Morin [26]’s integrative framework reaches similar results, spanning concepts of “reflective/extended” consciousness (higher self-reflection) and recursive self-awareness (i.e., awareness of being self-aware), buttressing the latter developed metacognitive knowledge. Although the entry points of the two frameworks differ, distinctions such as situational awareness and reflective self-awareness are consistently recognized.
Focusing on awareness, rather than consciousness, enables measurable, actionable progress in both cognitive science and AI, bridging conceptual divides and grounding research in functional, testable criteria.
a Metacognition
b Self-awareness
Figure 4: Illustration of metacognition and self-awareness as related but distinct components in awareness models
2.1 Major Types of Awareness
Metacognition
Metacognition, originally proposed as “thinking about thinking,” refers to the capacity to actively perceive, monitor, and regulate one’s own cognitive processes [3, 36, 37, 38, 39]. Nelson [36] distinguishes between metacognitive knowledge and metacognitive regulation, proposing a structural framework in which an object-level cognitive system provides input to a meta-level “central executive.” This central executive component monitors cognitive states through mechanisms such as confidence judgments (i.e., the association between task accuracy and confidence level [40]) and exerts control via strategic decisions and study-time allocation. Metacognitive knowledge encompasses a wide range of components: meta-level knowledge and beliefs pertain to an individual’s cognitive abilities, current tasks, past experiences, and specific process features (e.g., metamemory); metacognitive regulation involves active deployment of cognitive processes or resources, planning, monitoring, and strategic adjustments [41, 42, 43, 44, 45, 46, 47, 48]. During metacognitive regulation, an agent engages in continuous self-reflection and introspection, posing questions such as, “Am I likely to remember this information?” or “Will I deploy this module in the next operation?” and responds accordingly.
Extrapolating metacognitive processes to non-human agents remains controversial. Metacognition has traditionally been viewed as a uniquely human capacity [42, 39], with some scholars arguing that genuine metacognitive ability depends on linguistic structures that enable agents to attribute mental states to themselves [49]. Accumulating evidence suggests that certain non-human species, such as dolphins, primates, and birds, demonstrate behaviors indicative of meta-level cognitive processing [50, 51, 48]. For example, pigeons exhibit selective preferences for tasks requiring distinct working memory demands and engage in information-seeking behavior that mitigates the difficulty of discrimination tasks [52, 53]. Such evidence may suggest that pigeons monitor their knowledge states and thereby control their environment or adjust their problem-solving strategy. Nonetheless, without self-report instruments for animals, the evidence for animal metacognition remains contingent upon the interpretation of behavioral outcomes.
By analogy, AI agents endowed with metacognitive capabilities can perceive the expansion of their knowledge [54], assess confidence levels in their outputs [40], and adapt their reasoning strategies accordingly [48, 55]. Consider an AI-supported autonomous vehicle: its regulatory subsystem may supervise operational parameters and report errors, yet in the absence of agency or a self-reflective mechanism, such monitoring remains passive and reactive. It lacks the capacity to actively alter primary processes based on internal evaluation. In contrast, truly reflective behavior entails at least the capacity for self-monitoring—a hallmark of more advanced cognitive agents. Contemporary AI systems increasingly exhibit rudimentary forms of such metacognitive monitoring, including the ability to evaluate and revise their own cognitive operations [56, 7, 57].
@@ 2.1 Major Types of Awareness
In terms of behavioral capacity, Self-awareness represents the capacity of taking oneself as the object of awareness [58], yet it contains a collection of different self-oriented functions: agency, body ownership, self-recognization, interoception (representation of inner bodily state, such as hunger and pain), knowledge boundaries, and autobiographical memory [59, 60]. The self, as an apparatus that carries an individual’s subjective experience, operates with various levels of competence. As early as 1972, Duval and Wicklund [4] proposed that self-awareness arises when the agent’s attention is directed inward, contrasting with general environmental awareness. Later contributions from social-cognitive psychology frame self-awareness as an information-processing capability linked to self-schema (i.e., a cognitive framework about how individuals perceive, interpret, and behave in various situations) and mechanisms of self-regulation [26, 61, 62]. With the help of neuroimaging techniques, neuropsychology builds up sound self-awareness through lesion studies and cases of deficiency, such as dementia, Alzheimer’s disease, and anosognosia [^3] Meaning the lack of awareness of one’s own illness or deficits (Greek: a-, “without”, nosos, “disease”, gnosis, “knowledge”). Described by Joseph Babinski in 1914, it first characterized stroke patients with left paralysis who did not recognize their hemiplegia [63]. [64, 65, 24]. Based on these definitions, before claiming self-awareness, an individual should at least fuse sensory, proprioceptive, and cognitive data into a coherent agent identity and have access to declarative knowledge about self, stating “the body, the internal bodily state, the actions, the consequences of those actions, and those past memories belongs to me”.
Self-awareness is widely regarded as a hallmark of higher-order cognition [61]. By providing the information essential for metacognition, it is foundational for developing self-knowledge, facilitating introspection, enhancing emotional responses, and supporting adaptive self-control [31]. Some studies attribute self-awareness under the rubric of metacognition in the context of cognitive psychology [40], while Morin [26] recognized the differences between meta-self-awareness and perceptual-level self-awareness by extracting the conceptual information about oneself from perceptual information. For example, self-aware agents obtain the intuitive feeling of stomach pain and cramps after long-time starvation; after one’s attention shifts to the feeling of hunger, they create a reflexive meta-representational knowledge in their mind. In other words, the phenomenological content of self-awareness remains the discomfort in the stomach, not thoughts about feeling hungry. Neuroimaging reveals their distinctions as well: both are linked to the Default Mode Network (DMN) and its core regions; conscious experiences that are deemed essential for generating self-awareness persistently activate parallel limbic-network areas, specifically the medial prefrontal cortex/anterior cingulate cortex (ACC) and the precuneus/posterior cingulate cortex [31]. The neural substrates of metacognition are concentrated within frontal executive-function regions, e.g., the lateral frontopolar cortex (lFPC) and dorsal anterior cingulate cortex (dACC) play critical roles in monitoring decision uncertainty and adjusting strategies, suggesting that metacognition relies upon a distinct prefrontal system [66, 67].
All agents possess knowledge about themselves, but not all form a sufficient, structural knowledge system to support higher cognitive processes. Many animals can respond to inner stimuli or exhibit complex feedback behaviors, yet may lack the capacity to represent themselves as distinct entities or to generate self-referential content [68, 26]. Mirror self-recognition (MSR) has long been the classic test of self-awareness, and only some primates, elephants, and socially intelligent birds like magpies have been argued to succeed in the test [69, 70, 71]. Using MSR results as the single criterion is undoubtedly questionable; supportively, mammals and highly intelligent birds exhibit more features in autobiographical memories by matching the new environment with self-referential cues from past experiences [72, 73, 74]. In the context of artificial intelligence, it may be necessary to undertake a renewed frame of self-awareness, since AI systems display extraordinarily advanced capacities in certain dimensions (e.g., retrieving past environments, no matter in terms of accuracy, reproducibility, or velocity), while the implementation of a primitive sense of body ownership and agency in robots and of how the ontogenetic process shapes robotic self remains ambiguous [75]. Converging perspectives from psychology, neuroscience, and AI characteristics, self-awareness as an advanced cognitive feature may root in self-representation, embodiment, and other physical properties—not necessarily dependent on so-called “subjective qualia” [^4] Philosophical term for the mind-body problem, referring to introspectively accessible phenomenological aspects in some mental states, such as perceptual experiences, bodily sensations, moods, and emotional reactions [76]. [4, 62, 77].
a Social Awareness
b Situational Awareness
Figure 5: Illustration of social awareness and situational awareness as related but distinct components in awareness models
Social Awareness
Social Awareness is broadly defined as the cognitive capacity to perceive, interpret, and respond to the social signals, emotions, and perspectives of other agents [5]. This is a multifaceted construct encompassing theory of mind (ToM, i.e., the ability to attribute independent mental states such as beliefs, intentions, and knowledge to oneself other agents [78]), empathy, the understanding of interpersonal relationships, and the knowledge of society: context, cultural, and social norm (see 5a). Social awareness forms a foundational basis for self-construction within social contexts [61]. Individuals without neurological deviations gradually acquire the understanding that others possess autonomous beliefs and desires, along with the capacities for perspective-taking and affective empathy [78, 79, 80]. By approximately age four, typically developing children succeed in false-belief tasks, evidencing a functioning theory of mind [81], whereas children with autism spectrum disorder [^5] A neurodevelopmental disorder characterized by social communication and interaction deficits and repetitive motor behaviors [82]. frequently struggle with such tasks [83]. Humans further demonstrate exceptional proficiency in shared intentionality—the ability to collaboratively comprehend and align with others’ goals and perspectives [84].
Non-human species also exhibit foundational elements of social awareness. Primates [78] and birds [85] demonstrate rudimentary theory-of-mind capabilities, the cornerstone for extending emotional and relational knowledge. Animals with social structures and high cognitive functions exhibit pronounced forms of social awareness as well: chimpanzees and other primates can infer the goals and intentions of others and may even engage in deceptive behaviors [86]; corvids such as scrub-jays re-hide their food caches when previously observed, indicating awareness of potential pilferers [87]; dolphins recognize individual identities and maintain complex, multi-tiered alliances, suggesting an ability to attribute both knowledge and ignorance to conspecifics [88].
Early developments in artificial intelligence and robotics sought to model elementary components of social awareness [89, 90]. For instance, classical AI agents within multi-agent systems were designed to reason about the beliefs and intentions of other agents (e.g., [91]). Early social robotics integrated rudimentary theory-of-mind modeling and emotion-recognition mechanisms to support basic forms of human-robot interaction [92]. In AI contexts, social awareness entails perceiving and reasoning about the presence, internal states, and potential intentions of other agents (human or artificial). The criteria to identify competencies vary from recognition of social cues to more sophisticated forms of theory-of-mind tasks. For instance, a chatbot that detects user frustration from tone demonstrates external social sensitivity [93], whereas a robot that identifies informational gaps in its human collaborator and proactively offers relevant knowledge exemplifies a more advanced form of interpersonal reasoning [94, 95].